# Pretrained deep models outperform GBDTs in Learning-To-Rank under label scarcity

### v3

@@ 4.3 Peculiarities of best pretraining strategies for LTR
Our experiments suggest that the best pretraining and finetuning strategies for the LTR problem are different from the known best practices for image or text tasks. Particularly, we make two observations. (1) Full finetuning of the model performs much better than linear probing, which only finetunes a linear layer over frozen pretrained embeddings. (2) SimSiam and SimCLR-Rank perform significantly differently depending on the dataset. In the remainder of this section, we explore these two phenomena in more detail.
@@
Table 4: Finetuning: Comparison of fine tuning strategies on MSLR dataset (LP = Linear Probing, MP = Multilayer Probing, FF = Full Finetuning). Full finetuning consistently performs best or near-best, while linear probing and multilayer probing perform poorly .
@@
Figure 3: The question of whether SimCLR-Rank or SimSiam is better is dataset-dependent. Data points are averages over 3 trials. SimCLR-Rank performs better on MSLRWEB30K and Istella_S while SimSiam performs better on Yahoo Set1 under label scarcity.
Full finetuning outperforms linear probing. Linear probing is a popular finetuning strategy in text and image domains, where it produces good results (Chen & He, 2021; Chen et al., 2020; Peters et al., 2019; Kumar et al., 2022). We next investigate the performance of linear probing and full finetuning strategies in tabular LTR.
Results. Table 4 provides results on MSLR dataset. We find that in terms of NDCG of both SimCLR-Rank and SimSiam, LP is the worst and FF is the best. Interestly, MP comes as a close second in SimCLR-Rank, but it performs poorly on SimSiam. Our results on other datasets are similar (Section A.3.2). We thus recommend full fine-tuning as a stable and performant strategy.
In an attempt to explain the above observations, we plot t-SNE projections of the embeddings generated by the SimSiam and SimCLR-Rank encoders and the pre-final layer of a fully supervised DL model trained on all labels (Section A.3.3). These plots offer the following explanations for the two phenomena. (1) The embeddings produced by the SimSiam/SimCLR-Rank are not sorted by relevance score in the projection space, unlike the fully supervised encoder. This suggests that a linear ranker cannot use the pretrained embeddings to predict the true relevance labels. (2) MP works better for SimCLR-Rank than SimSiam because SimCLR-Rank’s embeddings are more evenly spread out (less collapsed) than those of SimSiam. This allows an MLP trained on the SimCLR-Rank embeddings to distinguish the items with different relevance labels.
SimCLR-Rank and SimSiam perform differently depending on the dataset. SimSiam and SimCLR-Rank perform significantly differently depending on the dataset, which is different from the observations of Chen & He (2021), who proposed SimSiam and found that it performs similarly to SimCLR derivatives. We next compare SimCLR-Rank with SimSiam at different levels of label scarcity.
(a) MSLR
(b) Yahoo Set1
Figure 4: t-SNE plots of SimSiam item embeddings for MSLR and Yahoo Set1 datasets. Relevance label is denoted by marker color/size. Compared to Yahoo Set1, MSLR has less spread out (more collapsed) embeddings and no clustering of relevant items.
Results. In Figure 3 we find that SimCLR-Rank performs better on MSLRWEB30K and Istella_S, while SimSiam performs much better on Yahoo Set1 when labels are scarce. Figure 4 shows the t-SNE projections of SimSiam item embeddings for MSLR and Yahoo Set1. MSLR embeddings are less spread out (more collapsed) and shows less desirable clustering of relevant items. We conjecture that this is caused by the high frequency of irrelevant items in MSLR.
@@ 5 Conclusion
6 Broader Impact
Our work focuses on improving the performing of tabular rankers. As stated in the introduction, tabular rankers are an instrumental component of many recommendation systems. These systems often use clicks as their learning targets, which can be problematic. For example, on social networks inflammatory and polarizing content may be more likely to garner clicks, which can have a negative societal impact. It is important to audit and responsibly deploy these recommendation systems, if they are used in areas that could potentially have large impacts on society.
7 Acknowledgements
@@ Appendix A Appendix
Figure 5: Simulated crowdsourcing of labels: We compare Outlier-NDCG (normal-$\uparrow$) pretrained rankers, non-pretrained DL rankers, and GBDT rankers as we change the percentage of training query groups that are labeled. To the left of the black dotted line, pretrained rankers perform the best. Points are averages over three trials. To the left of the black dotted vertical line, pretrained rankers are (1) significantly better on outliers than GBDTs at the $p=0.05$ level using a one-sided $t$-test, and (2) on average better on outliers than all other non-pretrained methods.
A.1 Detailed Related work
@@
Self-supervised learning Supervised learning of of
@@
MSLRWEB30K 18151 6072 Sub-sample pretrain SSL Val 6072 136 {0,1,2,3,4}
Yahoo Set1 14477 2147 Sub-sample pretrain SSL Val 5089 700 {0,1,2,3,4}
Istella_S 19200 7202 Sub-sample pretrain SSL Val 6523 220 {0,1,2,3,4}
@@
Self-supervised learning. Self-supervised learning (SSL) or unsupervised pretraining has improved performance in settings where there is a significant source of unlabeled data like text (Devlin et al., 2018) and images (Chen et al., 2020). In SSL, deep models are first pretrained on perturbed unlabeled data using self-supervised tasks to learn useful representations for the data. Then these models are finetuned for a downstream task with labeled data. Finetuning often takes one of two forms: (1) linear probing (a popular finetuning strategy in text and images (Chen & He, 2021; Chen et al., 2020; Peters et al., 2019)),where we freeze the pretrained model and only update the linear head during supervised finetuning, and (2) full finetuning, where we update the whole model during supervised finetuning (Devlin et al., 2018). Sometimes a mix of the two is used (Kumar et al., 2022). The core idea behind prominent SSL approaches like SimSiam and SimCLR is to carefully perturb input training samples, and train a representation that is consistent for transformations of the same sample. This provides robustness to natural perturbations and noise in data (Hendrycks et al., 2019).
Robustness in LTR. There has been prior work on studying worst-case behavior (robustness) of rankers (Voorhees, 2005; Zhang et al., 2013; Goren et al., 2018; Wu et al., 2022b, a; Penha et al., 2022). Some previous metrics measure a model’s robustness against adversarial attack (Goren et al., 2018; Wu et al., 2022a, 2023). Others measure the model’s per-query performance variance on a dataset (Voorhees, 2005; Zhang et al., 2013; Wu et al., 2022b). Our outlier metric, Outlier-NDCG, is a departure from previous work because it is not directly a measure of robustness–it is possible for a model to perform better on outlier data.
@@
Figure 6: An example of outlier detection in Istella (Lucchese et al., 2016) for our Outlier-NDCG metric.

### v4

@@ 4.3 Ablations: Choices for pretrained models in LTR
Our experiments suggest that the best pretraining and finetuning strategies for the LTR problem are different from the known best practices for image or text tasks. We make a few observations. (1) SimSiam and SimCLR-Rank perform significantly differently depending on the dataset. (2) Combining models pretrained by SimCLR-Rank and SimSiam can produce a model that is competitive against SimCLR-Rank and SimSiam across all datasets, producing a unified approach that may be performant across many settings. (3) Full finetuning of the model performs much better than linear probing, which only finetunes a linear layer over frozen pretrained embeddings. In the remainder of this section, we elaborate on these findings in more detail.
(1) SimCLR-Rank and SimSiam perform substantially differently depending on the dataset. SimSiam and SimCLR-Rank perform significantly differently depending on the dataset, which is different from the observations of Chen & He (2021), who proposed SimSiam and found that it performs similarly to SimCLR derivatives. We next compare SimCLR-Rank with SimSiam at different levels of label scarcity.
Figure 3: Comparison of SimCLR-Rank and SimSiam pretraining strategies on three public datasets with varying fraction of labeled query groups. Neither method consistently dominate the other in all the datasets. However, by combining SimCLR-Rank and SimSiam encoders, we can produce models that are competitive across all datasets. Data points are averages over 3 trials.
Figure 4: t-SNE plots of SimCLR-Rank, SimSiam, and fully supervised embeddings on Yahoo Set1 datasets. Relevance label is indicated by marker size/color. We see that SimCLR-Rank and SimSiam embeddings are qualitatively very different. Both SimCLR-Rank and SimSiam embeddings do not sort the data as well as fully supervised embeddings, which most effectively groups data points with relevance label 4 together.
Results. In Figure 3 we find that SimCLR-Rank performs better on MSLRWEB30K and Istella_S, while SimSiam performs much better on Yahoo Set1 when labels are scarce. Figure 4 shows the t-SNE projections of sub-sampled item embeddings learned by SimCLR-Rank, SimSiam, and pre-final layer of a fully supervised model for Yahoo Set1. We see that SimCLR-Rank and SimSiam also have qualitatively different embeddings: SimSiam’s embeddings are more clustered while SimCLR-Rank’s embeddings are less clustered. These observations raise the question of whether we can constructively combine them. We give similar t-SNE plots of other datasets in Section A.3.3
(2) Combining models pretrained by SimCLR-Rank and SimSiam produces a model that has the strengths of both. To make steps towards a single recommendation for unsupervised pretraining in LTR, we unify SimCLR-Rank and SimSiam into a single method “SimCLR-Rank + SimSiam”.
Methodology. To produce a “SimCLR-Rank + SimSiam” model, we finetune a linear layer over the embeddings from a finetuned SimCLR-Rank concatenated with the embeddings from a finetuned SimCLR-Rank model. More details about how we combine the models is given in Section A.3.7.
Results. In Figure 3 we find that “SimCLR-Rank + SimSiam” performs equal to or better than SimCLR-Rank and SimSiam across many different percentages of labeled query groups. While the combined model does not uniformly dominate across all labeled query group percentages, it is a promising first step towards a single unified method for unsupervised pretraining in ranking.
(3) Full finetuning outperforms linear probing. Linear probing is a popular finetuning strategy in text and image domains, where it produces good results (Chen & He, 2021; Chen et al., 2020; Peters et al., 2019; Kumar et al., 2022). We next compare it with the full finetuning strategy in tabular LTR.
Results. Table 4 provides results on MSLRWEB30K dataset. We find that in terms of NDCG of both SimCLR-Rank and SimSiam, LP is the worst and FF is the best. Interestingly, MP comes as a close second in SimCLR-Rank, but it performs poorly on SimSiam. Our results on other datasets are similar (Section A.3.2). We thus recommend full fine-tuning as a stable and performant strategy.
To explain the above observations, we we can look again at the t-SNE projections of the embeddings generated by the SimSiam and SimCLR-Rank encoders and the pre-final layer of a fully supervised DL model trained a randomly sampled 1000 samples from the full training set (Figure 4). These plots offer the following explanations for the two phenomena. (1) The embeddings produced by the SimSiam/SimCLR-Rank are not sorted by relevance score in the projection space, unlike the fully supervised encoder. This suggests that a linear ranker cannot use the pretrained embeddings to predict the true relevance labels. (2) MP works better for SimCLR-Rank than SimSiam because SimCLR-Rank’s embeddings are more evenly spread out (less collapsed) than those of SimSiam. This allows an MLP trained on the SimCLR-Rank embeddings to distinguish the items with different relevance labels.
@@
Table 4: Finetuning: Comparison of fine tuning strategies on MSLRWEB30K dataset (LP = Linear Probing, MP = Multilayer Probing, FF = Full Finetuning). Full finetuning consistently performs best or near-best, while linear probing and multilayer probing perform poorly.
@@ 5 Conclusion
Acknowledgements
@@ Appendix A Appendix
A.1 Detailed discussion of related work
@@
Self-supervised learning. Self-supervised learning (SSL) or unsupervised pretraining has improved performance in settings where there is a significant source of unlabeled data like text (Devlin et al., 2018) and images (Chen et al., 2020). In SSL, deep models are first pretrained on perturbed unlabeled data using self-supervised tasks to learn useful representations for the data. Then these models are finetuned for a downstream task with labeled data. Finetuning often takes one of two forms: (1) linear probing (a popular finetuning strategy in text and images (Chen & He, 2021; Chen et al., 2020; Peters et al., 2019)),where we freeze the pretrained model and only update the linear head during supervised finetuning, and (2) full finetuning, where we update the whole model during supervised finetuning (Devlin et al., 2018). Sometimes a mix of the two is used (Kumar et al., 2022). The core idea behind prominent SSL approaches like SimSiam and SimCLR is to carefully perturb input training samples, and train a representation that is consistent for perturbations of the same sample. This provides robustness to natural perturbations and noise in data (Hendrycks et al., 2019).
Robustness in LTR. There has been prior work on studying worst-case behavior (robustness) of rankers (Voorhees, 2005; Zhang et al., 2013; Goren et al., 2018; Wu et al., 2022b; a; Penha et al., 2022). Some previous metrics measure a model’s robustness against adversarial attack (Goren et al., 2018; Wu et al., 2022a; 2023). Others measure the model’s per-query performance variance on a dataset (Voorhees, 2005; Zhang et al., 2013; Wu et al., 2022b). Our outlier metric, Outlier-NDCG, is a departure from previous work because it is not directly a measure of robustness and it is possible for a model to perform better on outlier data.
A.2 Experimental Details
A.2.1 Dataset statistics
@@
Pretraining Supervised learning of of
@@
MSLRWEB30K 18151 6072 Subset Pretrain SSL Val 6072 136 {0,1,2,3,4}
Yahoo Set1 14477 2147 Subset Pretrain SSL Val 5089 700 {0,1,2,3,4}
Istella_S 19200 7202 Subset Pretrain SSL Val 6523 220 {0,1,2,3,4}
@@
A.2.2 Outliers query group details
Here we detail how we select the outlier query groups for the Outlier-NDCG metrics (Section 2.1). For our proprietary online shopping dataset, there are known outlier features that cause low quality predictions. These outliers arise from noise introduced in various stages in the data pipeline. We have mitigation strategies in place against these feature outliers, but they are often imperfect. Because we already know some of the outliers features, we deem any query group with items with outlier features as an outlier. These outliers comprise roughly 10% of the test set, which is still a large number of query groups (much larger than the number of query groups in the outlier datasets of the public datasets).
@@ A.2.2 Outliers query group details
Figure 5: An example of outlier detection in Istella_S (Lucchese et al., 2016) for our Outlier-NDCG metric.
Because different datasets have differently-sized typical gaps, we tune $G$ for each dataset (MSLR, Yahoo, Istella) such that the resulting percentage of outlier queries is as close to 1% of the test set as we can get. MSLRWEB30K has $G=5$, with 0.65% (40/6072) outlier queries, Yahoo has $G=20$, with 1.4% (30/2147) outlier queries, and Istella has $G=32$, with 0.46% (34/7202) outlier queries. 1% is a hyperparameter that can be tuned according to the user’s goals.
A.2.3 Model Settings and Training Details
Pretrained DL model.
Pretraining: (1) the encoder used for pretraining is the tabular ResNet (Gorishniy et al., 2021) with three ResNet blocks, with the final linear layer removed, (2) pretraining is done on the entire dataset with learning rate 0.0005 using Adam (Kingma & Ba, 2014) (3) with SimSiam/SimCLR-Rank/SimCLR-Sample, we tuned among four different augmentations for each pretraining method: randomly zeroing out features (“Zeroing”) with probabilities 0.1 or 0.7, and Gaussian noise with scale 1.0 or 2.0, (4) for SCARF we tune the corruption probability in $\{0.3,0.6,0.9\}$, (5) for VIME-self we tune the corruption probability in $\{0.3,0.5,0.7\}$, (6) for DACL+ we tune the mixup amount in $\{0.3,0.6,0.9\}$, (7) for SubTab we divide input features into 4 subsets with 75% overlap, and tune the masking probability in $\{0.3,0.6,0.9\}$, (8) we pretrain for 300 epochs for all methods, and (9) we use a batch size of roughly 200000 items for SimCLR-Rank, SimSiam, VIME-self, and SubTab as they are less GPU memory intensive while using a batch size of roughly 2000 items for SimCLR/DACL+/SCARF which are more GPU memory intensive.
Figure 6: We plot the t-SNE plots of embeddings produced by three different encoders for the MSLRWEB30K dataset: (1) pretrained by SimCLR-Rank, (2) pretrained by SimSiam, (3) trained via supervised training on the entire training set on roughly 1000 samples from MSLR. Marker size and color indicates relevance. We find (1) SimCLR-Rank/SimSiam cluster different relevances effectively but do not order them as well as the supervised encoder. (2) SimCLR-Rank produces more spread-out embeddings with less defined clusters than SimSiam.
No Pretrained DL model. There are a few rankers we evaluated for the no-pretrain DL baseline: (1) a 3-layer tabular ResNet from (Gorishniy et al., 2021) with a three-layer MLP on top of it, and was trained for 300 epochs on the labeled training set with learning rate tuned in {0.01, 0.001, 0.0001} using Adam, with early stopping using the validation set, (2) a DeepFM (Guo et al., 2017) model following the model architecture in the paper with learning rate tuned in {0.05, 0.005, 0.0005}, and (3) a DCNv2 (Wang et al., 2021) model with learning rate tuned in {0.05, 0.005, 0.0005} and cross layers tuned in {2,3,4}. DeepFM and DCNv2 were designed to use categorical features, and in MSLRWEB30K/Yahoo Set1/Istella_S do not document their categorical features (if any). We performed an analysis of the datasets and found 5, 48, 6 categorical features in MSLRWEB30K/Yahoo Set1/Istella_S respectfully each with cardinality at most 3. We embedded these categorical features into embeddings of dimension 8, following the rule-of-thumb in (Wang et al., 2021).
GBDT model. The GBDT ranker is the one from lightgbm (Ke et al., 2017) and we grid search the number of leaves in {31, 96, 200} minimum data in leaf in {20, 60, 200}, and column sample in {0.5, 0.9, 1.0} individually for each data point (for a total of 9 difference choices), while letting the rest of the parameters be the default in lightgbm (our tuning strategy is similar to Qin et al. (2021)). We use the LambdaRank loss to train the models.
GBDT + pseudolabeling. We add pseudo-labeled GBDTs as a semi-supervised GBDT baseline. For GBDTs with pseudo-labeling (GBDT-pseudo), we propose the following algorithm as there are no references to our knowledge. First, we train the GBDT on the labeled train set, and then use this GBDT to pseudo-labels for the unlabeled train set. A GBDT ranker outputs real-valued scores. To convert real-valued scores into relevance scores, we take the resulting scores and evenly bin them into 31 buckets. The scores converted to 0 are the lowest bucket, the scores converted to 1 are the second lowest bucket, and so on. We choose 31 because that is the largest number of labels allowed for GBDT ranking in lightgbm (the reason is because the LambdaRank loss requires the calculation of $2^{\textrm{label}}$ which can get very large).
GBDT + PCA. For GBDT+PCA, we take the entire training set’s features and learn a PCA on it with {5, 10, 15} components, similar to (Duh & Kirchhoff, 2008), to produce {5, 10, 15} new features (i.e. we tune among 5, 10, 15). We then train on the labeled train set, with both the original features and the new PCA features together.
GBDT + PCA + pseudolabeling. We also introduce a baseline where we perform PCA on the full training dataset, and then perform pseudolabeling as above. We tune the PCA similarly to the GBDT + PCA baseline.
All reported values for Outlier-NDCG are those achieved by the corresponding ranker with best overall NDCG. We report all results as averages over 3 trials, and use single V100 GPUs from a shared cluster to run the experiments.
A.2.4 Details of Binary Label generation used in Section 4.1.2
Following the methodology from Yang et al. (2022), we generate stochastic binary labels $y$ from the relevance labels $r$ as $y=\mathbbm{1}\{t\cdot r+G_{1}>t\cdot\tau_{\text{target}}+G_{0}\}$ where $t$ is a temperature parameter, $G_{1},G_{0}$ are standard Gumbels, and $\tau_{\text{target}}$ is a parameter controlling how sparse the binary labels are. Higher $\tau_{\text{target}}$ leads to higher label sparsity. Yang et al. (2022) show that $y$ is 1 with probability $\sigma(t\cdot(r-\tau_{\text{target}}))$ where $\sigma$ is the sigmoid function. We set $t=4$ as in Yang et al. (2022). For a given $\tau_{\text{target}}$, we produce produce a new dataset for each of MSLRWEB30K, Yahoo Set1, and Istella_S where we convert the relevance labels in the training/validation sets to binary labels and keep the true relevance labels in the test set.
A.2.5 Large industry-scale dataset experimental details
For pretraining, we used an augmentation which randomly zeroed out some features. We also tried randomly swapping some features between items in the same batch or query group, but they didn’t perform as well as zeros augmentation. During the fine-tuning stage we discard the projectors and predictor and add back the linear layer to complete the non-pretrained ranker structure. DL rankers are trained/fine-tuned using PiRank loss Swezey et al. (2021), batchsize of 1000, AdamW optimizer with learning rate of $0.00001$, weight decay of $0.01$ and cosine scheduling with warmup. All training stages and experiments were conducted on machines with 4 GPUs, where we ran 100k iterations with ddp and picked the best checkpoint based on validation metric.
A.3 Additional results
A.3.1 Runtime comparisons of pre-training strategies
Seconds per epoch
Method Complexity MSLRWEB30K Yahoo Set1 Istella
SimCLR $O(B^{2}L^{2})$ 102.81 $\pm$ 1.14 14.45 $\pm$ 0.04 69.26 $\pm$ 0.16
SimCLR-Rank (new) $O(BL^{2})$ 9.4700 $\pm$ 0.02 1.970 $\pm$ 0.01 5.950 $\pm$ 0.04
SimSiam $O(BL)$ 0.8700 $\pm$ 0.01 0.130 $\pm$ 0.00 0.730 $\pm$ 0.00
A.3.2 Additional Results on linear probing vs finetuning
Here we provide more results on the question of linear probing vs full finetuning, as discussed in Section 4.3. In Tables 7 and 8 we compare different probing strategies for all datasets when using SimSiam and SimCLR-Rank, respectively, as the pretraining method.
Table 7: SimSiam under different finetuning strategies on NDCG/Outlier-NDCG, averaged over 3 trials (LP = Linear Probing, MP = Multilayer Probing, FF = Full Finetuning). We find that linear probing and MLP probing perform extremely poorly (except in Yahoo Set1, where MLP probing performs well).
Method MSLRWEB30K ($\uparrow$) Yahoo Set1 ($\uparrow$) Istella ($\uparrow$)
NDCG ($\uparrow$)
LP 0.2679 $\pm$ 0.0007 0.6089 $\pm$ 0.0032 0.3805 $\pm$ 0.0042
MP 0.2764 $\pm$ 0.0001 0.6137 $\pm$ 0.0022 0.4484 $\pm$ 0.0020
FF 0.3935 $\pm$ 0.0034 0.6107 $\pm$ 0.0035 0.5618 $\pm$ 0.0049
Outlier-NDCG ($\uparrow$)
LP 0.1803 $\pm$ 0.0033 0.5088 $\pm$ 0.002 0.4407 $\pm$ 0.0388
MP 0.1749 $\pm$ 0.0023 0.5157 $\pm$ 0.0080 0.5324 $\pm$ 0.0002
FF 0.3149 $\pm$ 0.0119 0.52 $\pm$ 0.0133 0.6348 $\pm$ 0.0164
Table 8: SimCLR-Rank under different finetuning strategies on NDCG/Outlier-NDCG, averaged over 3 trials (LP = Linear Probing, MP = Multilayer Probing, FF = Full Finetuning). Multilayer probing performs moderately worse than full finetuning, and linear probing performs much worse than all other strategies.
Method MSLRWEB30K ($\uparrow$) Yahoo Set1 ($\uparrow$) Istella ($\uparrow$)
NDCG ($\uparrow$)
LP 0.3219 $\pm$ 0.0224 0.5202 $\pm$ 0.0093 0.4029 $\pm$ 0.0073
MP 0.3890 $\pm$ 0.0011 0.5942 $\pm$ 0.0016 0.5813 $\pm$ 0.0006
FF 0.3959 $\pm$ 0.0022 0.6022 $\pm$ 0.0013 0.5839 $\pm$ 0.0013
Outlier-NDCG ($\uparrow$)
LP 0.2304 $\pm$ 0.0332 0.3811 $\pm$ 0.0048 0.4717 $\pm$ 0.0166
MP 0.2969 $\pm$ 0.0009 0.4888 $\pm$ 0.0046 0.6369 $\pm$ 0.0045
FF 0.2892 $\pm$ 0.0025 0.5143 $\pm$ 0.0055 0.6352 $\pm$ 0.0140
A.3.3 Additional Results on SimCLR-Rank vs SimSiam
In this section we give additional t-SNE plots of the embeddings learned by SimCLR-Rank and SimSiam encoders, and the pre-final layer of the fully supervised ranker (trained with all labels) on MSLRWEB30K dataset (Figure 6), and Istella_S dataset (Figure 7). For the Yahoo Set1 dataset, SimSiam performs the best, but for the above two datasets SimCLR-Rank is the best.
A.3.4 Additional Results: Simulated crowdsourced label setting
Figure 8: Simulated crowdsourcing of labels: We compare Outlier-NDCG ($\uparrow$) pretrained rankers, non-pretrained DL rankers, and GBDT rankers as we change the percentage of training query groups that are labeled. To the left of the black dotted line, pretrained rankers perform the best. Points are averages over three trials. To the left of the black dotted vertical line, pretrained rankers are (1) significantly better on outliers than GBDTs at the $p=0.05$ level using a one-sided $t$-test, and (2) on average better on outliers than all other non-pretrained methods.
In this part of the appendix we provide additional results on binary label generation as detailed in Section 4.1.2. In Tables 9 and 10 we give the dataset statistics when using $\tau_{\text{target}}=4.25$ and $\tau_{\text{target}}=5.1$, respectively. For these two settings, the GBDT baseline does not contain semi-supervised runs as we found that semi-supervision tends to generally degrade GBDT performance; see Section A.4. The “No Pretrain” baseline uses the tabular ResNet architecture.
Dataset # query groups # labeled QGs # items per QG # positives per QG
train      val train      val train      val train      val
MSLRWEB30K 18151    6072 15.9%    16.2% 124.35    122.38 1.4%    1.4%
Yahoo Set1 14477    2147 11.1%    10.8% 30.02    30.32 6.2%    6.0%
Istella_S 19200    7202 49.0%    49.5% 106.36    94.98 1.4%    1.6%
Dataset # query groups # labeled QGs # items per QG # positives per QG
train      val train      val train      val train      val
MSLRWEB30K 18151    6072 1.1%    1.3% 124.35    122.38 0.8%    0.7%
Yahoo Set1 14477    2147 0.6%    0.6% 30.02    30.32 5.4%    5.1%
Istella_S 19200    7202 3.0%    2.7% 106.36    94.98 1.1%    1.3%